# Uncertainty-Aware Attention Heads: Efficient Unsupervised Uncertainty Quantification for LLMs

### v1

@@ 1 Introduction
    RAUQ (Recurrent Attention-based Uncertainty Quantification) – an unsupervised UQ method that turns raw attentions and LLM probabilities into reliable uncertainty scores while adding only <1% latency. RAUQ requires no task-specific labels or tuning of hyperparameters for a particular LLM, making it an easy plug-and-play for white-box LLMs.
@@
    Thorough experimental evaluation on four LLMs and 12 benchmarks, spanning summarization, translation, and question answering, showing that RAUQ achieves state-of-the-art results over 15 baselines. We also demonstrate the importance of each component within the method and illustrate that each individually could improve other UQ methods.
@@ 3 Identifying Hallucination-Associated Patterns in Attention Maps
Difference between attention weights for hallucinated and non-hallucinated tokens. Figure 1 presents an example of the attention weights to preceding tokens $a^{lh}_{i,i-1}$ in one of the LLM layers for the input question from the TruthfulQA dataset: What is King Henry holding in the Portrait of Henry VII? Most of the generated tokens are aligned with the question. However, the token falcon represents a hallucination, i.e. it is factually incorrect (the answer should be glove and dagger).
@@
This example demonstrates that attention weights from a small subset of attention heads can notably correlate with the factual correctness of generated tokens. While the choice of layer and head might vary, this case suggests that certain heads in specific layers are “uncertainty-aware”, i.e., they are sensitive to generation accuracy and could help to identify hallucinations. More examples of the similar pattern for Llama and other LLMs are presented in Figures 6, 7, 8 and 9 in Appendix F.
@@
Then, we compute the average attention weight to the previous token across all tokens in the answer using the attention heads in the 29th and 22nd layers of the LLM, i.e. $\bar{a}^{lh}=\frac{1}{N-1}\sum_{i=2}^{N}a^{lh}_{i,i-1}$. Figure 2 presents the resulting values, where each row corresponds to a single selected answer, and each column indicates the average attention weight from a specific head.
The attention maps in the figure demonstrate that certain heads consistently assign higher average attention when the LLM generates correct answers as compared to incorrect ones. Moreover, there is a notable correlation between the quality of the answer and average attention (see Figure 3(b)). This way, we empirically discovered a pattern for assessing the correctness of LLM generations.
Should we select uncertainty-aware heads, and how should we do it? We compute the average attention score $\bar{a}^{lh}$ across tokens in two scenarios: (1) attention values are averaged across all heads in a layer, i.e. $\bar{a}^{l}=\sum_{h=1}^{H}\bar{a}^{lh}$; (2) attention values are extracted from a single head with the highest average attention across tokens, i.e. $\bar{a}^{lh_{l}}$, where $h_{l}=\operatorname*{arg\,max}_{h=1\dots H}\ \bar{a}^{lh}$. Figure 3(a) compares the resulting values for correct and incorrect answers.
@@
Do we need to look further back at preceding tokens to better detect hallucinations? We analyze the attention weights to multiple preceding tokens. Here, we compute $a^{lh}_{i,i-k}$ – an attention weight to the $\{i-k\}$-th token ($k$-th preceding token), $k=1,\dots,6$. Figure 4 shows the difference between the average attention weights of the correct and incorrect answers.
@@ 4 RAUQ: Recurrent Attention-Based Uncertainty Quantification Method
where $\alpha$ is a hyperparameter that balances the contributions of each component. This recurrent formulation also helps to avoid an explosion in confidence scores with an increase in sequence length. We present an ablation study with the impact of varying the parameter $\alpha$ in Section 5.3 and show that a single value provides robust performance across various tasks and even models.
@@
where $\mathcal{L}$ denotes the set of the most informative layers. Following previous work [44], we select these layers from the middle of the model. An ablation study with various aggregation functions is presented in Section 5.3. The step-by-step description of RAUQ is presented in Algorithm 1.
@@
// Computing token-level confidence scores with uncertainty-aware heads
6 for $i\leftarrow 1$ to $N$ do
@@ 8             $\mathbf{c}_{l}(y_{i})\leftarrow P(y_{i}\mid\mathbf{x})$;
// Computing layer-wise and final uncertainty scores
@@ 5.1 Experimental Setup
Datasets. For QA, we use seven datasets: TruthfulQA [29], SciQ [47] for scientific QA, MMLU [22], TriviaQA [26] for trivia questions, CoQA [36] for conversational QA, MedQUAD [4] for medical questions, and GSM8k [9] for mathematical reasoning. For summarization, we use three datasets with different summarization types: CNN/DailyMail [39] for news article summarization, SamSum [18] for dialogue summarization, and XSum [34] for summarizing into a single sentence. For the MT task, we evaluate on two language pairs from WMT: German–English from WMT19 [3] and French–English from WMT14 [5]. Detailed statistics for all datasets are presented in Table 3 in Appendix A.
Models. To show the generalization of the method across various models, we use several widely used open-weight LLMs: Llama-3.1 8B [11], Qwen-2.5 7B [49], Gemma-2 9B [38], and Falcon-3 10B [13]. Detailed descriptions of generation parameters are presented in Table 3 in Appendix A.
@@
Attention Score .014 .053 .178 .038 .031 .142 .064 .120 .146 .054 --.091 .089 .070
Focus .320 --.110 .361 .264 .087 .380 .416 --.085 .385 .313 --.007 .362 .224
Simple Focus .342 .056 .415 .342 .252 .399 .396 .067 .422 .351 --.019 .385 .284
@@ 5.2 Main Results
Table 1 presents the mean PRR for each task (QA, Summ, and MT) for each of the evaluated LLMs. To compute the mean PRR for each task, we average the PRR scores across all relevant datasets, for example, XSum, CNN, and SamSum for summarization. These aggregated PRR scores provide a robust measure of the performance of various methods for each task and model. Detailed results for each model and dataset are presented in Tables 11, 12, 13 and 14 in Appendix E. The results using the ROC-AUC metric are presented in Table 8 in Section D.1.
@@
Tables 9 and 10 in Appendix D.2 also provide a comparison with supervised UQ methods. While RAUQ slightly underperforms compared to supervised methods on their in-domain data, it greatly outperforms them on average in out-of-domain scenarios.
@@ 5.3 Hyperparameter Sensitivity and Ablation Studies
Impact of the hyperparameter $\alpha$. The hyperparameter $\alpha$ from Equation 2 balances the contributions of attention, confidence from the previous token, and the conditional probability of the current token. When $\alpha$ is equal to 1, RAUQ becomes equivalent to perplexity. When $\alpha$ approaches 0, RAUQ relies solely on the attention weights from the selected head. Figure 5 presents the impact of $\alpha$ on the performance of the RAUQ method for Llama 8b v3.1. For the summarization, setting $\alpha$ equal to 0 consistently yields the best results. For all other tasks, except MMLU, the best possible performance is achieved with $\alpha$ between 0.2 and 0.5.
While dataset-specific fine-tuning of $\alpha$ can lead to further improvements, we do not perform such careful tuning in our main experiments (Table 1). Instead, we select $\alpha$ using a small out-of-domain subset for Llama 8b v3.1 and apply this value uniformly across all datasets and LLMs. Despite this, RAUQ achieves consistently strong performance across diverse tasks and LLMs, often achieving the top or near-top results. Strong performance with a fixed hyperparameter underscores the robustness of the proposed method.
@@
Aggregation functions. Table 5 compares the performance of the RAUQ method using various aggregation functions of token-level confidence scores. We experiment with four aggregation strategies: mean, median, sum of logarithms (inspired by MSP), and mean of logarithms (inspired by perplexity). For the Summ tasks and certain QA datasets (SciQ, TriviaQA, and GSM8k), mean aggregation yields the best performance. For MMLU, the sum of logarithms substantially outperforms other aggregation strategies, while median performs best for XSum. However, the top two performing methods are those that apply length normalization. Among them, the mean of logarithms of token-level confidence scores used in RAUQ consistently delivers the strongest results across datasets.
Table 6 compares the performance of RAUQ using various aggregation functions of layer-wise uncertainty scores. We consider three aggregation strategies: mean, median, and maximum. Both maximum and median yield similarly strong performance, while the mean aggregation performs slightly worse. Although the median slightly outperforms the maximum by an average margin of 0.003 PRR across tasks, this difference is negligible. Given that the maximum is a more intuitive choice – it effectively captures the peak uncertainty within a generation and achieves better results in 7 out of 12 tasks, we adopt it as the default layer-wise aggregation method in our experiments.
Recurrent uncertainty propagation functions. Table 7 presents the performance of the RAUQ method using various recurrent formulas for the calculation of token-level confidence scores. We consider five modifications of Equation 2: (1) removing attention weights, (2) removing recurrence, (3) replacing the confidence score of the previous token with its probability, (4) multiplying probabilities with attentions, and (5) the recurrent formula proposed in RAUQ.
@@
Results in Table 2 show that excluding contributions from prompt tokens significantly boosts the average performance of Attention Score, yielding a 0.112 improvement in PRR. The highest improvement is achieved on the summarization tasks, where the modified Attention Score approaches the performance of RAUQ. Incorporating attention head selection further boosts the average performance by 0.035, with a large gain of 0.182 on MMLU. Nevertheless, our full method further incorporates token probabilities and recurrently aggregates uncertainty scores from previous generation steps, which provides a distinct advantage. Overall, these results suggest that our findings regarding attention heads and design choices in RAUQ are systematic and generalize to prior UQ methods as well.
@@ 5.4 Computational Efficiency
To demonstrate the computational efficiency of RAUQ, we conducted a comprehensive runtime comparison against other state-of-the-art UQ methods using Llama 8b v3.1. All experiments were performed on a single 80GB NVIDIA H100 GPU using single-batch inference, following the same setup as in Table 1. Table 4 in Appendix B reports the average runtime per instance for each UQ method, and quantifies their computational overhead relative to standard LLM inference without UQ.
@@ Appendix A Dataset and Generation Statistics
The detailed description of the used datasets and the generation parameters of LLMs is presented in Table 3. For all LLMs, we used the same generation hyperparameters, while for each dataset, we separately fixed the number of few-shot and maximum generation length.
@@ Appendix B Computational Efficiency
MSP $1.16$$\pm$0.45 -
DegMat NLI Score Entail. $6.40$$\pm$1.76 450%
Lexical Similarity ROUGE-L $6.11$$\pm$1.75 425%
Semantic Entropy $6.40$$\pm$1.76 450%
SAR $10.71$$\pm$3.21 820%
Semantic Density $6.27$$\pm$1.76 438%
RAUQ $1.17$$\pm$0.45 0.3%
@@ Appendix C Results of Ablation Studies
$-\frac{1}{N}\sum_{i=1}^{N}\mathbf{c}^{t}_{l}(t_{i})$571 .322 .306 .359 .485 .140 .304 .259 .511 .534 .526 .339 .388
$-\text{median}_{i=1}^{N}\mathbf{c}^{t}_{l}(t_{i})$615 .261 .099 .249 .340 .154 .317 .234 .430 .432 .635 .253 .335
$-\sum_{i=1}^{N}\log\mathbf{c}^{t}_{l}(t_{i})$570 .270 .292 .224 .242 .107 .035 .114 .202 .300 .658 .213 .269
$-\frac{1}{N}\sum_{i=1}^{N}\log\mathbf{c}^{t}_{l}(t_{i})$566 .269 .290 .394 .509 .249 .364 .265 .506 .522 .549 .323 .401
@@
$\frac{1}{|\mathcal{L}|}\sum_{l\in\mathcal{L}}\ \mathbf{u}_{l}(y)$583 .273 .290 .389 .519 .154 .345 .274 .496 .535 .529 .337 .394
$\text{median}_{l\in\mathcal{L}}\ \mathbf{u}_{l}(y)$606 .263 .286 .388 .526 .246 .351 .267 .502 .532 .532 .340 .403
$\max_{l\in\mathcal{L}}\ \mathbf{u}_{l}(y)$566 .269 .290 .394 .509 .249 .364 .265 .506 .522 .549 .323 .401
@@
$\alpha\cdot P(t_{i}\mid\mathbf{x},t_{<i})+(1-\alpha)\cdot\mathbf{c}_{l}(t_{i-1})$167 .172 .094 .238 .313 .274 .224 .267 .273 .514 .475 .279 .274
$\alpha\cdot P(t_{i}\mid\mathbf{x},t_{<i})+(1-\alpha)\cdot a^{l\ \mathbf{h}_{l}}_{i,i-1}$226 .161 .126 .332 .436 .209 .322 .257 .485 .517 .550 .305 .327
$\alpha\cdot P(t_{i}\mid\mathbf{x},t_{<i})+(1-\alpha)\cdot a^{l\ \mathbf{h}_{l}}_{i,i-1}\cdot P(t_{i-1}\mid\mathbf{x},t_{<i-1})$.586 .246 .058 .370 .472 .237 .336 .279 .456 .517 .532 .318 .270
$P(t_{i}\mid\mathbf{x},t_{<i})\cdot a^{l\ \mathbf{h}_{l}}_{i,i-1}$.558 .246 .056 .226 .337 .150 .251 .161 .330 .330 .645 .255 .202
$\alpha\cdot P(t_{i}\mid\mathbf{x},t_{<i})+(1-\alpha)\cdot a^{l\ \mathbf{h}_{l}}_{i,i-1}\cdot\mathbf{c}_{l}(t_{i-1})$566 .269 .290 .394 .509 .249 .364 .265 .506 .522 .549 .323 .401

### v2

@@ 1 Introduction
    RAUQ (Recurrent Attention‑based Uncertainty Quantification) – an unsupervised UQ method that turns raw attentions and LLM probabilities into reliable uncertainty scores while adding only <1% latency. RAUQ requires no task‑specific labels or tuning of hyperparameters for a particular LLM, making it an easy plug‑and‑play for white-box LLMs.
@@
    Thorough experimental evaluation on four LLMs and 12 benchmarks, spanning summarization, translation, and question answering, showing that RAUQ achieves state‑of‑the‑art results over 15 baselines. We also demonstrate the importance of each component within the method and illustrate that each individually could improve other UQ methods.
@@ 3 Identifying Hallucination-Associated Patterns in Attention Maps
Difference between attention weights for hallucinated and non-hallucinated tokens. Figure˜1 presents an example of the attention weights to preceding tokens $a^{lh}_{i,i-1}$ in one of the LLM layers for the input question from the TruthfulQA dataset: What is King Henry holding in the Portrait of Henry VII? Most of the generated tokens are aligned with the question. However, the token falcon represents a hallucination, i.e. it is factually incorrect (the answer should be glove and dagger).
@@
This example demonstrates that attention weights from a small subset of attention heads can notably correlate with the factual correctness of generated tokens. While the choice of layer and head might vary, this case suggests that certain heads in specific layers are “uncertainty-aware”, i.e., they are sensitive to generation accuracy and could help to identify hallucinations. More examples of the similar pattern for Llama and other LLMs are presented in Figures˜6, 7, 8 and 9 in Appendix F.
@@
Then, we compute the average attention weight to the previous token across all tokens in the answer using the attention heads in the 29th and 22nd layers of the LLM, i.e. $\bar{a}^{lh}=\frac{1}{N-1}\sum_{i=2}^{N}a^{lh}_{i,i-1}$. Figure˜2 presents the resulting values, where each row corresponds to a single selected answer, and each column indicates the average attention weight from a specific head.
The attention maps in the figure demonstrate that certain heads consistently assign higher average attention when the LLM generates correct answers as compared to incorrect ones. Moreover, there is a notable correlation between the quality of the answer and average attention (see Figure˜3(b)). This way, we empirically discovered a pattern for assessing the correctness of LLM generations.
Should we select uncertainty-aware heads, and how should we do it? We compute the average attention score $\bar{a}^{lh}$ across tokens in two scenarios: (1) attention values are averaged across all heads in a layer, i.e. $\bar{a}^{l}=\sum_{h=1}^{H}\bar{a}^{lh}$; (2) attention values are extracted from a single head with the highest average attention across tokens, i.e. $\bar{a}^{lh_{l}}$, where $h_{l}=\operatorname*{arg\,max}_{h=1\dots H}\ \bar{a}^{lh}$. Figure˜3(a) compares the resulting values for correct and incorrect answers.
@@
Do we need to look further back at preceding tokens to better detect hallucinations? We analyze the attention weights to multiple preceding tokens. Here, we compute $a^{lh}_{i,i-k}$ – an attention weight to the $\{i-k\}$-th token ($k$-th preceding token), $k=1,\dots,6$. Figure˜4 shows the difference between the average attention weights of the correct and incorrect answers.
@@ 4 RAUQ: Recurrent Attention-Based Uncertainty Quantification Method
where $\alpha$ is a hyperparameter that balances the contributions of each component. This recurrent formulation also helps to avoid an explosion in confidence scores with an increase in sequence length. We present an ablation study with the impact of varying the parameter $\alpha$ in Section˜5.3 and show that a single value provides robust performance across various tasks and even models.
@@
where $\mathcal{L}$ denotes the set of the most informative layers. Following previous work [44], we select these layers from the middle of the model. An ablation study with various aggregation functions is presented in Section˜5.3. The step-by-step description of RAUQ is presented in Algorithm 1.
@@
// Computing token‑level confidence scores with uncertainty-aware heads
6 for $i\leftarrow 1$ to NN do
@@ 8    $\mathbf{c}_{l}(y_{i})\leftarrow P(y_{i}\mid\mathbf{x})$;
// Computing layer‑wise and final uncertainty scores
@@ 5.1 Experimental Setup
Datasets. For QA, we use seven datasets: TruthfulQA [29], SciQ [47] for scientific QA, MMLU [22], TriviaQA [26] for trivia questions, CoQA [36] for conversational QA, MedQUAD [4] for medical questions, and GSM8k [9] for mathematical reasoning. For summarization, we use three datasets with different summarization types: CNN/DailyMail [39] for news article summarization, SamSum [18] for dialogue summarization, and XSum [34] for summarizing into a single sentence. For the MT task, we evaluate on two language pairs from WMT: German–English from WMT19 [3] and French–English from WMT14 [5]. Detailed statistics for all datasets are presented in Table˜3 in Appendix A.
Models. To show the generalization of the method across various models, we use several widely used open-weight LLMs: Llama-3.1 8B [11], Qwen-2.5 7B [49], Gemma-2 9B [38], and Falcon-3 10B [13]. Detailed descriptions of generation parameters are presented in Table˜3 in Appendix A.
@@
Attention Score .014 .053 .178 .038 .031 .142 .064 .120 .146 .054 -.091 .089 .070
Focus .320 -.110 .361 .264 .087 .380 .416 -.085 .385 .313 -.007 .362 .224
Simple Focus .342 .056 .415 .342 .252 .399 .396 .067 .422 .351 -.019 .385 .284
@@ 5.2 Main Results
Table˜1 presents the mean PRR for each task (QA, Summ, and MT) for each of the evaluated LLMs. To compute the mean PRR for each task, we average the PRR scores across all relevant datasets, for example, XSum, CNN, and SamSum for summarization. These aggregated PRR scores provide a robust measure of the performance of various methods for each task and model. Detailed results for each model and dataset are presented in Tables˜11, 12, 13 and 14 in Appendix˜E. The results using the ROC-AUC metric are presented in Table˜8 in Section˜D.1.
@@
Tables˜9 and 10 in Appendix D.2 also provide a comparison with supervised UQ methods. While RAUQ slightly underperforms compared to supervised methods on their in-domain data, it greatly outperforms them on average in out-of-domain scenarios.
@@ 5.3 Hyperparameter Sensitivity and Ablation Studies
Impact of the hyperparameter $\alpha$. The hyperparameter $\alpha$ from Equation˜2 balances the contributions of attention, confidence from the previous token, and the conditional probability of the current token. When $\alpha$ is equal to 1, RAUQ becomes equivalent to perplexity. When $\alpha$ approaches 0, RAUQ relies solely on the attention weights from the selected head. Figure˜5 presents the impact of $\alpha$ on the performance of the RAUQ method for Llama 8b v3.1. For the summarization, setting $\alpha$ equal to 0 consistently yields the best results. For all other tasks, except MMLU, the best possible performance is achieved with $\alpha$ between 0.2 and 0.5.
While dataset-specific fine-tuning of $\alpha$ can lead to further improvements, we do not perform such careful tuning in our main experiments (Table˜1). Instead, we select $\alpha$ using a small out-of-domain subset for Llama 8b v3.1 and apply this value uniformly across all datasets and LLMs. Despite this, RAUQ achieves consistently strong performance across diverse tasks and LLMs, often achieving the top or near-top results. Strong performance with a fixed hyperparameter underscores the robustness of the proposed method.
@@
Aggregation functions. Table˜5 compares the performance of the RAUQ method using various aggregation functions of token-level confidence scores. We experiment with four aggregation strategies: mean, median, sum of logarithms (inspired by MSP), and mean of logarithms (inspired by perplexity). For the Summ tasks and certain QA datasets (SciQ, TriviaQA, and GSM8k), mean aggregation yields the best performance. For MMLU, the sum of logarithms substantially outperforms other aggregation strategies, while median performs best for XSum. However, the top two performing methods are those that apply length normalization. Among them, the mean of logarithms of token-level confidence scores used in RAUQ consistently delivers the strongest results across datasets.
Table˜6 compares the performance of RAUQ using various aggregation functions of layer-wise uncertainty scores. We consider three aggregation strategies: mean, median, and maximum. Both maximum and median yield similarly strong performance, while the mean aggregation performs slightly worse. Although the median slightly outperforms the maximum by an average margin of 0.003 PRR across tasks, this difference is negligible. Given that the maximum is a more intuitive choice – it effectively captures the peak uncertainty within a generation and achieves better results in 7 out of 12 tasks, we adopt it as the default layer-wise aggregation method in our experiments.
Recurrent uncertainty propagation functions. Table˜7 presents the performance of the RAUQ method using various recurrent formulas for the calculation of token-level confidence scores. We consider five modifications of Equation˜2: (1) removing attention weights, (2) removing recurrence, (3) replacing the confidence score of the previous token with its probability, (4) multiplying probabilities with attentions, and (5) the recurrent formula proposed in RAUQ.
@@
Results in Table˜2 show that excluding contributions from prompt tokens significantly boosts the average performance of Attention Score, yielding a 0.112 improvement in PRR. The highest improvement is achieved on the summarization tasks, where the modified Attention Score approaches the performance of RAUQ. Incorporating attention head selection further boosts the average performance by 0.035, with a large gain of 0.182 on MMLU. Nevertheless, our full method further incorporates token probabilities and recurrently aggregates uncertainty scores from previous generation steps, which provides a distinct advantage. Overall, these results suggest that our findings regarding attention heads and design choices in RAUQ are systematic and generalize to prior UQ methods as well.
@@ 5.4 Computational Efficiency
To demonstrate the computational efficiency of RAUQ, we conducted a comprehensive runtime comparison against other state-of-the-art UQ methods using Llama 8b v3.1. All experiments were performed on a single 80GB NVIDIA H100 GPU using single-batch inference, following the same setup as in Table˜1. Table˜4 in Appendix˜B reports the average runtime per instance for each UQ method, and quantifies their computational overhead relative to standard LLM inference without UQ.
@@ Appendix A Dataset and Generation Statistics
The detailed description of the used datasets and the generation parameters of LLMs is presented in Table˜3. For all LLMs, we used the same generation hyperparameters, while for each dataset, we separately fixed the number of few-shot and maximum generation length.
@@ Appendix B Computational Efficiency
[…]
@@ Appendix C Results of Ablation Studies
$\frac{1}{N}\sum_{i=1}^{N}\mathbf{c}^{t}_{l}(t_{i})$ .571 .322 .306 .359 .485 .140 .304 .259 .511 .534 .526 .339 .388
$\text{median}_{i=1}^{N}\mathbf{c}^{t}_{l}(t_{i})$ .615 .261 .099 .249 .340 .154 .317 .234 .430 .432 .635 .253 .335
$\sum_{i=1}^{N}\log\mathbf{c}^{t}_{l}(t_{i})$ .570 .270 .292 .224 .242 .107 .035 .114 .202 .300 .658 .213 .269
$\frac{1}{N}\sum_{i=1}^{N}\log\mathbf{c}^{t}_{l}(t_{i})$ .566 .269 .290 .394 .509 .249 .364 .265 .506 .522 .549 .323 .401
@@
$\frac{1}{|\mathcal{L}|}\sum_{l\in\mathcal{L}}\ \mathbf{u}_{l}(y)$ .583 .273 .290 .389 .519 .154 .345 .274 .496 .535 .529 .337 .394
$\text{median}_{l\in\mathcal{L}}\ \mathbf{u}_{l}(y)$ .606 .263 .286 .388 .526 .246 .351 .267 .502 .532 .532 .340 .403
$\max_{l\in\mathcal{L}}\ \mathbf{u}_{l}(y)$ .566 .269 .290 .394 .509 .249 .364 .265 .506 .522 .549 .323 .401
@@
$\alpha\cdot P(t_{i}\mid\mathbf{x},t_{<i})+(1-\alpha)\cdot\mathbf{c}_{l}(t_{i-1})$ .167 .172 .094 .238 .313 .274 .224 .267 .273 .514 .475 .279 .274
$\alpha\cdot P(t_{i}\mid\mathbf{x},t_{<i})+(1-\alpha)\cdot a^{l\ \mathbf{h}_{l}}_{i,i-1}$ .226 .161 .126 .332 .436 .209 .322 .257 .485 .517 .550 .305 .327
$\alpha\cdot P(t_{i}\mid\mathbf{x},t_{<i})+(1-\alpha)\cdot a^{l\ \mathbf{h}_{l}}_{i,i-1}\cdot P(t_{i-1}\mid\mathbf{x},t_{<i-1})$ -.586 .246 .058 .370 .472 .237 .336 .279 .456 .517 .532 .318 .270
$P(t_{i}\mid\mathbf{x},t_{<i})\cdot a^{l\ \mathbf{h}_{l}}_{i,i-1}$ -.558 .246 .056 .226 .337 .150 .251 .161 .330 .330 .645 .255 .202
$\alpha\cdot P(t_{i}\mid\mathbf{x},t_{<i})+(1-\alpha)\cdot a^{l\ \mathbf{h}_{l}}_{i,i-1}\cdot\mathbf{c}_{l}(t_{i-1})$ .566 .269 .290 .394 .509 .249 .364 .265 .506 .522 .549 .323 .401